### v3

# A Graph-Based Classical and Quantum Approach to Deterministic L-System Inference

Abstract.
L-systems can be made to model and create simulations of many biological processes, such as plant development. Finding an L-system for a given process is typically solved by hand, by experts, in a massively time-consuming process. It would be significant if this could be done automatically from data, such as from sequences of images. In this paper, we are interested in inferring a particular type of L-system, deterministic context-free L-system (D0L-system) from a sequence of strings. We introduce the characteristic graph of a sequence of strings, which we then utilize to translate our problem (inferring D0L-systems) in polynomial time into the maximum independent set problem (MIS) and the SAT problem. After that, we offer a classical exact algorithm and an approximate quantum algorithm for the problem.
1. Introduction
Lindenmayer systems, abbreviated as L-systems, were originally introduced in [15]. They are rewriting systems or formal grammar systems where the letters of a string are rewritten in parallel, and this process continues to produce a sequence of strings. One of their applications is modeling plant growth and biological systems, such as plant morphology [22]. When the letters of an L-system are interpreted as drawing instructions, L-system simulators such as vlab [26], can produce accurate graphical depictions of plant development over time.
Constructing L-systems algorithmically from a sequence of images could make L-systems and their creation more practical. This is because developing an L-system for a growing plant currently requires expertise in L-system construction and is extremely time-consuming [6]. The process of deriving an L-system from data is known as the inductive inference problem. The automated process of inference of L-systems falls into an area of machine learning. Solving this problem would allow for the automatic creation of models and simulations. The images that are produced by these simulations could be used in different ways, for example to create synthetic data to enrich training sets for supervised machine learning applications, which typically lack diversity [7]. Inductive inference is the problem of taking one or more sequences of strings generated by an unknown L-system, and to infer an L-system that initially generates these strings.
While classical algorithms have been proposed for the inductive inference problem for deterministic L-systems (D0L-systems, [2]) and stochastic L-systems ([3], [17]), there is still room for improvement through faster and more efficient algorithms, approximate methods, or quantum algorithms. For D0L-systems, it is known that the problem can be solved in polynomial time if the size of the alphabet is fixed [18], but it is NP-complete if the alphabet is not fixed [9]. Quantum computing has shown promising results in approaching machine learning problems ([27, 28, 16, 14, 21, 23, 12, 4]) and NP-complete problems ([10, 19, 1]). The main motivation of this paper is to build a bridge between quantum computing and the inductive inference problem.
This paper is a step toward developing classical and approximate quantum algorithms for the D0L inductive inference problem (inferring a D0L-system from a sequence of strings, hereafter D0LII). To achieve this, we present a theorem that relates D0LII to the maximum independent set (MIS) problem. Specifically, we propose a polynomial time and size encoding of inputs of D0LII into a graph, which we call the characteristic graph. We then show that the D0LII solution is equivalent to finding an MIS of a specific size in the characteristic graph. This encoding and the related theorem allow us to propose different solutions to D0LII. One solution is a quantum algorithm that solves the MIS through the Quantum Approximate Optimization Algorithm (QAOA, [10, 25, 5]) and then infers production rules. Another solution is an exact algorithm, a combination of an exact solution for MIS, followed by inferring the D0L-system rules. A third solution, which we will not discuss in detail but is evident, involves using a polynomial encoding of MIS into SAT and then employing modern SAT solvers. We will also examine advancements in quantum computing, particularly concerning MIS and how it could be modified to target the version of MIS to which we are concerned.
This paper is organized as follows: Section 2 provides the necessary background on L-systems, the MIS problem, and the notations required for the presented quantum algorithm. Section 3 presents polynomial encoding into MIS and SAT. Section 4 reviews the QAOA hybrid quantum algorithm for MIS. In the final section, we propose quantum algorithms for D0LII.
2. Preliminaries and Notations
This section introduces the foundational concepts and notations required for our work. First, we define 0L-systems and D0L-systems, formally specify the MIS problem, and introduce quantum notations relevant to our approach. These preliminaries are needed to encode the D0LII into MIS.
In the study of formal language theory, an alphabet $V$ is a finite set of symbols. Furthermore, $V^{*}$ represents the set of all strings over $V$, which includes the empty word $\varepsilon$. A language over $V$ is any subset of $V^{*}$. Given any string $S$, we use the notation $|S|$ to refer to its length and $S[i]$ to refer to the $i$-th ($1\leq i\leq|S|$) character of $S$. Furthermore, for $i$ and $j$ where $11\leq i\leq j\leq|S|+$ 1, we define $S[i:j]$ as follows:
  • If $i=j$, then it is the empty string.
  • If $i<j<|S|+1$, then it is the substring from index $i$ (inclusive) to $j$ (exclusive).
  • If $i<j=|S|+1$, then it is the suffix starting from index $i$ (inclusive).
Please see [13] for an introduction to formal language theory and computational complexity. We also refer to the cardinality of a finite set $X$ as simply $|X|$, which represents the number of elements in $X$.
2.1. 0L-systems and D0L-systems
Next, we introduce L-systems and related notations, following the approaches of [18] and [22]. A context-free L-system, also known as a 0L-system, applies productions to symbols independently of their context within a string. We denote an L-System as $G=(V,\omega,P)$, where $V$ is an alphabet, $\omega\in V^{*}$ is the axiom, and $P\subseteq V\times V^{*}$ is a finite set of productions. In the literature, L-systems that allow productions to the empty word are sometimes called propagating. We write a production $(a,x)\in P$ as $a\rightarrow x$, where $a$ is called the predecessor, and $x$ is called the successor.
Typically, we assume that each $a\in V$ has at least one production $a\rightarrow x$ in $P$. However, for mathematical convenience, we may consider “partial” L-systems where this property does not hold. A 0L-system $G$ is deterministic (and called a D0L-system) if each $a\in V$ has exactly one production with it as a predecessor. For a string $\mu=a_{1}\cdots a_{n}\in V^{*}$, we write $\mu\Rightarrow\nu$ to indicate that $\mu$ directly derives $\nu$ if $\nu=x_{1}\cdots x_{n}$, where $a_{i}\rightarrow x_{i}\in P$ for all $1\leq i\leq n$. A derivation $d$ in $G$ comprises:
  1. (1)
    A trace $(w_{0},\ldots,w_{m})$ where $w_{i}\Rightarrow w_{i+1}$ for $0\leq i<m$.
  2. (2)
    A function $\sigma_{d}:\{(i,j)\mid 0\leq i<m,1\leq j\leq|w_{i}|\}\rightarrow P$ such that if $w_{i}=A_{1}\cdots A_{|w_{i}|}$ and $1\leq j\leq|w_{i}|$, then $w_{i+1}=\alpha_{1}\ldots\alpha_{|w_{i}|}$ where $\sigma_{d}(i,j)=A_{j}\rightarrow\alpha_{j}$.
The function $\sigma_{d}$ specifies which productions are applied to each letter in the derivation.
A sequence $\theta=(w_{0},\ldots,w_{m})$ is said to be compatible with a 0L system $G$ if $G$ can generate $\theta$ as a trace; otherwise, it’s incompatible. With these definitions, we can now formally state D0LII:
{questionbox}
D0L inductive inference problem or D0LII: Given alphabet $V$, and a sequence $\theta=(w_{0},\ldots,w_{m})$ over $V$, construct a D0L $G=(V,\omega,P)$ such that $\theta$ is compatible with $G$, or decide that is not possible?
2.2. Maximum Independent Set (MIS) Problem
In this part, we will define the MIS problem for graphs. In an undirected graph $G=(V(G),E(G))$, a set $I\subseteq V(G)$ is called an independent set, if the induced graph on $I$ does not contain any edges; that is, there is no edge between any $u,v\in V(I)$. An independent set $I$ is termed a maximum independent set if it satisfies the following condition:
$|I|=\max\left\{|J|:J\subseteq V(G),J\text{ is independent}\right\}.$
The problem of finding a MIS in a graph is NP-complete ([11]).
2.3. Quantum Notations
In the sections where we discuss the quantum algorithm for MIS, we will use the following notations, which are adopted from [20]:
  • $Z_{i}$: Pauli-Z operator on qubit $i$
  • $X_{i}$: Pauli-X operator on qubit $i$
  • $R_{Z_{i}}(\theta)$: Rotation around Z-axis by angle $\theta$ on qubit $i$
  • $R_{X_{i}}(\theta)$: Rotation around X-axis by angle $\theta$ on qubit $i$
  • |$\ket{+}=\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$: The superposition state
  • $\delta_{ij}$: Kronecker delta, defined as $\delta_{ij}=\begin{cases}1&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$
  • $I$ and $\bar{1}$ are the identity matrix and a vector with all components being $1$ respectively.
Having proposed the main question we intend to solve, we are now ready to encode the D0LII into MIS.
3. Encoding D0LII into MIS and SAT
In this section, we will discuss how to map an instance of D0LII into MIS and SAT in polynomial time. One way to find an encoding from D0LII to SAT is to establish an encoding from D0LII to MIS. Since a polynomial reduction exists from MIS to SAT, denoted as $f$, it is sufficient to construct a polynomial reduction, such as $g$, from the D0LII to MIS. The reduction from MIS to SAT is known ([11], and is implied from their NP-completeness). This relationship can be visualized as in Figure 1.
[…]
Figure 1. Illustration of the relationship between D0LII, MIS, and SAT.
In order to find the encoding from D0LII to MIS, we will introduce the characteristic graph of a D0LII problem which will be a simple undirected graph. Such a graph encodes the information of $\theta$ and the restriction of D0LII into a single graph. Given the sequence $\theta=(w_{0},\ldots,w_{m})$, let $G_{\theta}$ be the characteristic graph of a D0LII with the following vertex set:
  • For $10\leq i\leq m-$ 1 and $1\leq j\leq|w_{i}|$, we construct complete graphs $G_{i,j}$ such that:
    • If $j=1=|w_{i}|$:
      $V(G_{i,j})=\{(i,j,1,|w_{i+1}|+1)\};$
    • If $j=1$ and $j<|w_{i}|$:
      $V(G_{i,j})=\{(i,j,1,\text{end}):1\leq\text{end}\leq|w_{i+1}|+1\};$
    • If $1<j$ and $j=|w_{i}|$:
      $V(G_{i,j})=\{(i,j,\text{start},|w_{i+1}|+1):1\leq\text{start}\leq|w_{i+1}|+1\};$
    • If $1<j$ and $j<|w_{i}|$:
      $V(G_{i,j})=\{(i,j,\text{start},\text{end}):1\leq\text{start}\leq\text{end}\leq|w_{i+1}|+1\}.$
  • We then define:
    $V(G_{\theta})=\bigcup\limits_{0\leq i<m,1\leq j\leq|w_{i}|}V(G_{ij}).$
Hereafter, we define $E(G_{\theta})$, which includes all edges in
$\bigcup\limits_{0\leq i<m,\ 1\leq j\leq|w_{i}|}E(G_{ij}).$
Also let $(i_{1},j_{1})\neq(i_{2},j_{2})$ and $v,w\in V(G_{\theta})$ such that
$v\in V(G_{i_{1},j_{1}})\quad\text{and}\quad w\in V(G_{i_{2},j_{2}}).$
Then we write
$v=(i_{1},j_{1},s_{1},e_{1})\quad\text{and}\quad w=(i_{2},j_{2},s_{2},e_{2}).$
The edge $(v,w)\in E(G_{\theta})$ if at least one of the following conditions holds:
  1. (1)
    $w_{i_{1}}[j_{1}]=w_{i_{2}}[j_{2}]$ and $w_{i_{1}+1}[s_{1}:e_{1}]\neq w_{i_{2}+1}[s_{2}:e_{2}]$;
  2. (2)
    $i_{1}=i_{2}$, $j_{2}=j_{1}+1$ and $e_{1}\neq s_{2}$.
Having defined $G_{\theta}$, we are ready to present our theorem relating D0LII to MIS.
Theorem 3.1.
Assume $V=\{a_{1},\ldots,a_{n}\}$ is an alphabet, and $\theta=(w_{0},\ldots,w_{m})$ is a sequence over $V$, then there exists a D0L-system $G$ such that $\theta$ is compatible with $G$ if and only if $G_{\theta}$ has a MIS of size $k$, where:
$k=\sum_{i=0}^{m-1}\lvert w_{i}\rvert.$
Proof.
()($\Rightarrow$)( ⇒ ): Suppose D0L-system $G$ with production set $P$ derives $\theta$. We will show that $G_{\theta}$ has a MIS of size $k$. Since $G_{\theta}$ induced on $G_{i,j}$ is a complete graph, then any independent set has size at most $k$. Therefore, it remains to be shown that some independent set attains this upper bound. $G$ derives $\theta$, so there exists a derivation $d$ in $G$ consisting of:
  1. (1)
    A trace $(w_{0},\ldots,w_{m})$ where $w_{i}\Rightarrow w_{i+1}$ for $0\leq i<m$.
  2. (2)
    A function $\sigma_{d}:\{(i,j)\mid 0\leq i<m,1\leq j\leq|w_{i}|\}\rightarrow P$ such that if $w_{i}=A_{1}\cdots A_{|w_{i}|}$ and $1\leq j\leq|w_{i}|$, then $w_{i+1}=\alpha_{1}\ldots\alpha_{|w_{i}|}$ where $\sigma_{d}(i,j)=A_{j}\rightarrow\alpha_{j}$.
Let $Z_{i,j}=|\alpha_{j}|$, where $\sigma_{d}(i,j)=\alpha_{j}$ for $0\leq i<m$, $1\leq j\leq|w_{i}|$, and define:
[…]
Certainly, $I=k$. It remains to show that $I$ is an independent set. Assume the contrary, there exists $(i_{1},j_{1},s_{1},e_{1}),(i_{2},j_{2},s_{2},e_{2})\in I$ such that:
$\left((i_{1},j_{1},s_{1},e_{1}),(i_{2},j_{2},s_{2},e_{2})\right)\in E(G_{\theta}).$
For such an edge to exist, the pair must satisfy at least one of the conditions for edges of $G_{\theta}$. This edge can not exist due to the first condition since if:
$w_{i_{1}}[j_{1}]=w_{i_{2}}[j_{2}],$
then having that $P$ is the production set of a D0L-system and the way $I$ is constructed, we get:
$w_{i_{1}+1}[s_{1}:e_{1}]=w_{i_{2}+1}[s_{2}:e_{2}],$
which would give us that:
$\left((i_{1},j_{1},s_{1},e_{1}),(i_{2},j_{2},s_{2},e_{2})\right)\notin E(G_{\theta}),$
which is a contradiction. Therefore, we assume such an edge exists due to the second condition. Then we have one of the following two cases:
  1. (1)
    $i_{1}=i_{2}$ and $j_{2}=j_{1}+1$ and $e_{1}\neq s_{2}$;
  2. (2)
    $i_{1}=i_{2}$ and $j_{1}=j_{2}+1$ and $e_{2}\neq s_{1}$.
The first case has contradiction with $G$ deriving $\theta$, since from $w_{i}$ to $w_{i+1}$, the image of $w_{i_{1}}[j_{1}]$ is not coming right before the image of $w_{i_{2}}[j_{2}]$. Similarly the second case leads to contradiction as well. Therefore, neither of the conditions could be satisfied for
$\left((i_{1},j_{1},s_{1},e_{1}),(i_{2},j_{2},s_{2},e_{2})\right)$
to be in $E(G_{\theta})$ which is a contradiction. Therefore, $I$ is a max independent set of size $k$.
()($\Leftarrow$)( ⇐ ): For the backwards direction we assume that $G_{\theta}$ has MIS $I$ of size $k$. We define the axiom of our D0L-system $G$ to be $\omega=w_{0}$. Since the induced subgraph to $G_{i,j}$ is complete, then from each $G_{i,j}$ there exists a unique vertex in $I$. We identify such vertex with $I_{i,j}$. It is left to build the production set $P$. Define:
$P=\{w_{i}[j]\rightarrow w_{i+1}[s_{i,j}:e_{i,j}]\mid 0\leq i<m,1\leq j\leq\lvert w_{i}\rvert,(i,j,s_{i,j},e_{i,j})=I_{i,j}\}.$
This production set, along with $\omega$, will give us an 0L-system. We first show that this 0L-system is a D0L-system. Assume the contrary: there exist $I_{i_{1},j_{1}},I_{i_{2},j_{2}}\in I$ such that
$I_{i_{1},j_{1}}=(i_{1},j_{1},s_{i_{1},j_{1}},e_{i_{1},j_{1}})\quad\text{and}\quad I_{i_{2},j_{2}}=(i_{2},j_{2},s_{i_{2},j_{2}},e_{i_{2},j_{2}}),$
with the following conditions:
$w_{i_{1}}[j_{1}]\to w_{i_{1}+1}[s_{i_{1},j_{1}}:e_{i_{1},j_{1}}]\quad\text{and}\quad w_{i_{2}}[j_{2}]\to w_{i_{2}+1}[s_{i_{2},j_{2}}:e_{i_{2},j_{2}}]\in P,$
such that
$w_{i_{1}+1}[s_{i_{1},j_{1}}:e_{i_{1},j_{1}}]\neq w_{i_{2}+1}[s_{i_{2},j_{2}}:e_{i_{2},j_{2}}],$
and
$w_{i_{1}}[j_{1}]=w_{i_{2}}[j_{2}].$
However, by the first condition of edges, we have:
$(I_{i_{1},j_{1}},I_{i_{2},j_{2}})\in E(G_{\theta}),$
which is a contradiction with $I$ being an independent set. Therefore, the 0L-system $G$ is a D0L-system. It remains to show that $G=(V,\omega,P)$ derives $\theta$. More specifically, we need to show that for $0\leq i<m$, $G$ applied to $w_{i}$ will give us $w_{i+1}$. $G$ applied to $w_{i}$, will give us the following:
$w_{i+1}[s_{i,1}:e_{i,1}]w_{i+1}[s_{i,2}:e_{i,2}]\ldots w_{i+1}[s_{i,|w_{i}|}:e_{i,|w_{i}|}],$
where the last string is the concatenation of the substrings $w_{i+1}[s_{i,j}:e_{i,j}]$ for $1\leq j\leq|w_{i+1}|$. From the construction of $V_{G_{i,1}}$ and $V_{G_{i,|w_{i}|}}$, we know that:
$s_{i,1}=1\text{ and }e_{i,|w_{i}|}=|w_{i+1}|+1.$
Furthermore, since $I_{i,j}$ are vertices of an independent set, and by the second condition of edges $G_{\theta}$, we get that for $1\leq j<|w_{i}|$ we have:
$e_{i,j}=s_{i,j+1},$
finishing the proof that $G$ from $w_{i}$ derives $w_{i+1}$. Since $i$, $0\leq i<m$ was arbitrary, $G$ derives $\theta$ finishing the proof of the backward direction.
Corollary 3.2.
Given $\theta=(w_{0},\ldots,w_{m})$, if $G_{\theta}$ has a MIS $I$ of size $k$ (where $k$ is as defined in the statement of the theorem) in $G_{\theta}$, then the D0L-system $(V,w_{0},P)$ generates $\theta$ where:
  1. (1)
    $V=\{w_{i}[j]\mid 0\leq i<m,1\leq j\leq|w_{i}|\}$,
  2. (2)
    $P=\{w_{i}[j]\rightarrow w_{i+1}[s_{i,j}:e_{i,j}]\mid 0\leq i<m,1\leq j\leq|w_{i}|,(i,j,s_{i,j},e_{i,j})\in I\}$.
Corollary 3.3.
The D0LII can be polynomially encoded into an MIS by constructing $G_{\theta}$, and the production rules can be extracted from the vertices selected by the algorithm solving the MIS. By utilizing the polynomial encoding of MIS to SAT, the D0LII can be polynomially encoded into SAT.
4. Quantum Algorithm for the Modified MIS
There have been recent developments in quantum computing regarding the MIS problem, specifically, the introduction of a quantum polynomial approximate algorithm by [29] and a quantum algorithm with time complexity $\mathcal{O}(1.1488^{n})$ [8]. The latter offers Grover-based algorithms for both maximal independent set and maximum independent set problems. However, since the oracle in the algorithm was not clearly specified, we developed our own approach more similar to [29]. The approach we will practice here for solving MIS is based on turning the MIS into a QUBO (Quadratic Unconstrained Binary Optimization) matrix and then utilizing a QAOA to find an approximate solution. To see this, first note that for a graph $G$ with $n$ vertices, any candidate for an independent set for this graph can be represented by a vector of size $n$ such as $x=(x_{1},\ldots,x_{n})$ where $x_{i}\in\{0,1\}$ for $1\leq i\leq n$. Furthermore, let $Q$ be an $n$ by $n$ matrix such that:
  • $Q_{i,j}=-1$ if $i=j$;
  • $Q_{i,j}=1$ if $(i,j)\in E(G)$;
  • $Q_{i,j}=0$ if $(i,j)\notin E(G)$.
Then, we see in the following proposition that an MIS induces a vector $x$ that minimizes $x^{\top}Qx$, where $x^{\top}$ is the transpose of the column vector $x$.
Proposition 4.1.
Given a graph $G=(V(G),E(G))$ with $V(G)=\{v_{1},\ldots,v_{n}\}$ and a matrix $Q$ as the corresponding QUBO matrix, an MIS of $G$ will minimize the function $f(x)=x^{\top}Qx$.
Proof.
First, notice that:
$f(x)=\sum_{i}Q_{i,i}x_{i}+\sum_{(i,j)\in E(G)}Q_{i,j}x_{i}x_{j}.$
Given $T\subset V(G)$, let $x_{T}$ denote the corresponding binary vector for $T$.
Claim 4.2.
For any set $T$ that is not independent, there exists $S\subset T$ such that $f(x_{S})<f(x_{T})$.
Proof of Claim.  Let $T$ be a non-independent set, and take $a,b\in V(T)$ such that $(a,b)\in E(T)$. Define $S$ as the induced graph on $V(T)\setminus\{b\}$. Then we have:
$f(x_{T})=-|V(T)|+2|E(T)|,$
and
$f(x_{S})$ $=-|V(S)|+2|E(S)|$
$=-(|V(T)|-1)+2|E(S)|$
$\leq-(|V(T)|-1)+2(|E(T)|-1)$
$=-|V(T)|+1+2|E(T)|-2$
$=-|V(T)|+2|E(T)|-1$
$<f(x_{T}).$
Claim 4.3.
For any set $T$ that is not independent, there exists $S\subset T$ such that $S$ is independent and $f(x_{S})<f(x_{T})$.
Proof of Claim.  At each step, pick two vertices $a,b\in V(T)$ such that $(a,b)\in E(T)$. Apply Claim 4.2 to obtain a smaller subgraph with fewer edges such that $f$ has a smaller value on the subgraph. Repeat this process until arriving at an independent set $S$, where $f(x_{S})<f(x_{T})$, as desired.
Claim 4.4.
Given two independent sets $M$ and $N$ such that $|V(N)|<|V(M)|$, we have:
$f(x_{M})<f(x_{N}).$
Proof of Claim.  The proof follows directly from observing:
$f(x_{M})$ $=-|V(M)|+2|E(M)|$
$=-|V(M)|+0$
$<-|V(N)|$
$=-|V(N)|+0$
$=-|V(N)|+2|E(N)|$
$<f(x_{N}).$
Finally, to prove the proposition, let $M$ be any MIS of $G$ and $T$ any subset of $G$. If $T$ is independent, the claim is done by Claim 4.4. Otherwise, assume $T$ is not independent. By Claim 4.3, we can find an independent set $N\subset T$ such that:
(4.1) $f(x_{N})<f(x_{T}).$
Furthermore, since $M$ is an MIS and $N$ is an independent set, we have $|V(N)|\leq|V(M)|$. By Claim 4.4, it follows that:
(4.2) $f(x_{M})<f(x_{N}).$
Combining Equations 4.1 and 4.2, we conclude:
$f(x_{M})<f(x_{T}),$
completing the proof. ∎
Furthermore, if one is given that the graph given has an MIS of size at most $k$, and then looking for such MIS, then one could add the term $(x^{\top}\mathbf{1}-k)^{2}$ to the cost function. This will make the cost function as follows and would lead to a modification of the MIS quantum solver. Notice that in the following equations, since we are looking to minimize the cost function, then we can drop the constant terms which are indepedent of $x$. Therefore, in such cases, we use $\equiv$, and in cases where we have an actual equality, we will use $=$.
$cost(x)$ $=$ $x^{\top}Qx+\lambda(x^{\top}\bar{1}-k)^{2}$
$=$ $x^{\top}Qx+\lambda(\sum_{i=1}^{n}x_{i})^{2}-2\lambda k(\sum_{i=1}^{n}x_{i})+\lambda k^{2}$
$\equiv$ $x^{\top}Qx+\lambda(\sum_{i=1}^{n}x_{i})^{2}-2\lambda k(\sum_{i=1}^{n}x_{i})$
$=$ $x^{\top}Qx+\lambda x^{\top}Ix-2\lambda k(\sum_{i=1}^{n}x_{i})$
$=$ $x^{\top}(Q+\lambda I)x-2\lambda k(\sum_{i=1}^{n}x_{i})$
$=$ $\sum_{i,j}(Q_{ij}+\lambda\delta_{ij})x_{i}x_{j}-\lambda k\sum_{i}x_{i}.$
In the above equations, since we care about solutions which minimize $H_{C}$, we can drop the $\lambda k^{2}$ term. Furthermore, $\lambda$ is a parameter which we’ll use to control how much weight should be on the condition of the independent set being as close to $k$ as possible. Therefore, overall, we will have:
$cost(x)\equiv\sum_{i,j}(Q_{ij}+\lambda\delta_{ij})x_{i}x_{j}-\lambda k\sum_{i}x_{i}.$
We then map our classical QUBO to a quantum Hamiltonian. This is done by replacing each binary variable $x_{i}$ with the quantum operator $\frac{1}{2}(I-Z_{i})$, where $Z_{i}$ is the Pauli-Z operator on qubit $i$. The problem Hamiltonian is given by:
$H_{C}=\frac{1}{4}\sum_{i,j}(Q_{ij}+\lambda\delta_{ij})(I-Z_{i})(I-Z_{j})-\lambda k\sum_{i}(I-Z_{i}).$
Expanding this and using the fact that $Z_{i}Z_{j}=Z_{j}Z_{i}$, we obtain:
$H_{C}=\text{constant}+\sum_{i,j}\frac{1}{4}(Q_{ij}+\lambda\delta_{ij})Z_{i}Z_{j}-\sum_{i}\left(\frac{1}{2}\sum_{j}(Q_{ij}+\lambda\delta_{ij})-\lambda k\right)Z_{i}.$
The constant term can be ignored, as it does not affect the optimization. The next step is to define the mixing Hamiltonian for QAOA, which is:
$H_{M}=\sum_{i=1}^{n}X_{i},$
where $X_{i}$ is the Pauli-X operator on qubit $i$. Therefore, overall we have our cost and mixing Hamiltonians which are:
(4.3) $H_{C}=\sum_{i,j}\frac{1}{4}(Q_{ij}+\lambda\delta_{ij})Z_{i}Z_{j}-\sum_{i}\left(\frac{1}{2}\sum_{j}(Q_{ij}+\lambda\delta_{ij})-\lambda k\right)Z_{i},$
(4.4) $H_{M}=\sum_{i=1}^{n}X_{i}.$
Then one layer of the QAOA circuit models the following two unitary operators which are:
  • $e^{-i\gamma H_{C}}=\prod_{i,j}R_{Z_{i}Z_{j}}(\frac{\gamma}{4}(Q_{ij}+\lambda\delta_{ij}))\prod_{i=1}^{n}R_{Z_{i}}(\sum_{j}(Q_{ij}+\lambda\delta_{ij})-\lambda k)$,
  • $e^{-i\beta H_{M}}=\prod_{i=1}^{n}R_{X_{i}}(2\beta)$.
Then for a p-layer QAOA, we have parameters $\gamma_{1},\beta_{1},\ldots,\gamma_{p},\beta_{p}$, along with $\lambda$. Note that we fix $\lambda$ at the beginning of the problem, and update $\gamma_{i}$ and $\beta_{i}$. Therefore, our circuit $circ$ be equivalent to the following unitary transformation:
$circ=\prod_{i=1}^{p}(e^{-i\gamma_{i}H_{C}}e^{-i\beta_{i}H_{M}}).$
Algorithm 1 presents the complete modified algorithm for solving the modified version of the MIS problem for graphs $G$ where we know the MIS of $G$ has a size of at most $k$. Overall, this is a hybrid algorithm. We update the parameters using the subgradient method, and each iteration works as follows: First, we initialize the parameters $\bar{\gamma}$ and $\bar{\beta}$. Then, we enter a for loop where, for each iteration, the quantum circuit $circ$ is run $e$ times. The average of the $cost(x)$ is calculated, contributing to the overall average cost. Finally, we use a classical algorithm, such as the subgradient method, to update $\bar{\gamma}$ and $\bar{\beta}$ based on this average cost. Finally, the best $x$, which minimizes the cost corresponding to the MIS with the largest possible size, is returned.
Algorithm 1 QAOA for MIS
1:procedure Modified.QAOA.MIS.Solver($G,k,p$)
2:     Form matrix $Q$ from $G$
3:     Form $H_{C}$ and $H_{M}$ from Equations 4.3 and 4.4 respectively
4:     Initialize $\lambda$ to a large value
5:     Initialize $\vec{\gamma}=(\gamma_{1},\ldots,\gamma_{p})$ and $\vec{\beta}=(\beta_{1},\ldots,\beta_{p})$ randomly
6:     Set learning rate $\eta$ and number of iterations $T$
[…]
9:     for $t=1$ to $T$ do
10:         $\text{cost}_{\text{avg}}\leftarrow 0$
[…]
12:         for $i=1$ to $e$ do
13:              Prepare initial state |$\ket{+}^{\otimes n}$
14:              for $l=1$ to $p$ do
15:                  Apply $e^{-i\gamma_{l}H_{C}}$
16:                  Apply $e^{-i\beta_{l}H_{M}}$
17:              end for
18:              Measure the state to obtain bit string $x$
19:              Calculate $\text{cost}(x)=x^{\top}(Q+\lambda I)x-2\lambda k(\sum_{i=1}^{n}x_{i})$
20:              $\text{cost}_{\text{avg}}\leftarrow\text{cost}_{\text{avg}}+\text{cost}(x)$
21:              if $(x)<\text{best\_cost}$_cost then
[…]
24:              end if
25:         end for
26:         $\text{cost}_{\text{avg}}\leftarrow\text{cost}_{\text{avg}}/\text{e}$
27:         Calculate subgradient $\nabla\text{cost}_{\text{avg}}$ with respect to $\vec{\gamma}$ and $\vec{\beta}$
28:         Update $\vec{\gamma}\leftarrow\vec{\gamma}-\eta\nabla_{\vec{\gamma}}\text{cost}_{\text{avg}}$
29:         Update $\vec{\beta}\leftarrow\vec{\beta}-\eta\nabla_{\vec{\beta}}\text{cost}_{\text{avg}}$
30:     end for
31:     return best_set
32:end procedure
5. Quantum and Classical Algorithms for D0LII
Having Theorem 3.1 and Corollary 3.2, we are now ready to present our QAOA hybrid and classical Algorithms, which are Algorithms 2 and 3 respectively. Both algorithms start by constructing the character set $V$, which will be the characters appearing in the first $m$ strings of the sequence (all but the last). The axiom of our D0L-system should be the first element of the sequence. It then initializes the production set $P$ to the empty set. We then take $\theta$ to be the input sequence $(w_{0},\ldots,w_{m})$ and construct $G_{\theta}$. Algorithm 2 provides an approximate solution to the MIS (based on Algorithm 1), while Algorithm 3 finds the MIS of $G_{\theta}$ in an exact way. Algorithm 3 then calculates $k$. By Theorem 3.1, if $|I|<k$, the answer to the D0LII is negative, returning false. Otherwise, similar to the backward direction of the proof of Theorem 3.1, we can construct the L-system. Having $P$, we will return the D0L-system $(V,\omega_{0},P)$ as our D0LII. Notice that Algorithm 2 is not doing the check $|I|<k$ since this algorithm is an approximate algorithm and is not intended to find the exact solution to D0LII.
Algorithm 2 Solving D0LII Using QAOA
1:procedure QuantInferD0L($w_{0},w_{1},\dots,w_{m},p$)
2:     $V\leftarrow\{w_{i}[j]\mid 0\leq i<m,1\leq j\leq|w_{i}|\}$
3:     $\omega\leftarrow w_{0}$
4:     $P\leftarrow\{\}$
5:     $\theta\leftarrow(w_{0},w_{1},\dots,w_{m})$
6:     construct $G_{\theta}$
7:     $k\leftarrow\sum_{i=0}^{m-1}\lvert w_{i}\rvert$
8:In the following function call, for higher accuracy we need to increase $p$
9:     $I\leftarrow\text{Modified.QAOA.MIS.Solver}(G_{\theta},k,p)$
10:     for $(i,j,s,e)\in I$ do
11:         add $w_{i}[j]\rightarrow w_{i+1}[s:e]$ to $P$
12:     end for
13:     return $(V,\omega,P)$
14:end procedure
Algorithm 3 Solving D0LII Using an MIS Classical Solver
1:procedure ClassicalD0LSolver($w_{0},w_{1},\dots,w_{m}$)
2:     $V\leftarrow\{w_{i}[j]\mid 0\leq i<m,1\leq j\leq|w_{i}|\}$
3:     $\omega\leftarrow w_{0}$
4:     $P\leftarrow\{\}$
5:     $\theta\leftarrow(w_{0},w_{1},\dots,w_{m})$
6:     construct $G_{\theta}$
7:     $I\leftarrow MIS_{solver}(G_{\theta})$
8:     $k\leftarrow\sum_{i=0}^{m-1}\lvert w_{i}\rvert$
9:     if $|I|<k$ then
10:         return $False$
11:     end if
12:     for $(i,j,s,e)\in I$ do
13:         add $w_{i}[j]\rightarrow w_{i+1}[s:e]$ to $P$
14:     end for
15:     return $(V,\omega,P)$
16:end procedure
Similar to our algorithm, one could use the polynomial encoding of MIS to SAT and solve the D0LII problem using either quantum SAT solvers or classical SAT solvers. We will omit the details of this algorithm here since the idea is very similar to Algorithm 2.
Notice that the time complexity of our algorithms in the classical and quantum cases is dominated by solving the MIS. Before discussing time complexity, we define:
$l=\max_{0\leq i<m}|w_{i+1}|,$
and let $h$ be the frequency of the most frequently occurring character in $\theta$, $v$ to be the size of the alphabet ($v$ is the number of distinct characters appearing in $\theta$ other than the last string), and recall that $k=\sum_{i=0}^{m-1}|w_{i}|$. To find the time complexity of the classical MIS solver, we start by noting the following:
$|V(G_{\theta})|$ $=$ $\sum_{i=0}^{m-1}\sum_{j=1}^{|w_{i}|}|V(G_{i,j})|$
$\leq$ $\sum_{i=0}^{m-1}\sum_{j=1}^{|w_{i}|}l^{2}$
$=$ $\sum_{i=0}^{m-1}|w_{i}|l^{2}$
$=$ $l^{2}\sum_{i=0}^{m-1}|w_{i}|$
$=$ $kl^{2}.$
Therefore, $|V(G_{\theta})|=\mathcal{O}(kl^{2})$, and given the best-known classical MIS solver [24], the worst-case time complexity of the classical MIS solver is $\mathcal{O}(2^{|V(G_{\theta})|/4})=\mathcal{O}(2^{kl^{2}/4})$. Therefore, the worst-case time complexity of Algorithm 3 is also $\mathcal{O}(2^{kl^{2}/4})$.
The time complexity of Algorithm 2 is dominated by MIS quantum solver, which involves the following steps:
  • Constructing the QUBO Matrix $Q$, with time complexity $\mathcal{O}((kl^{2})^{2})$.
  • Quantum circuit preparation and state preparation: Initializing the quantum register in the superposition state |$\ket{+}^{\otimes\mathcal{O}(kl^{2})}$, with time complexity $\mathcal{O}(kl^{2})$.
  • Cost Hamiltonian $H_{C}$: Applying a rotation gate $R_{Z_{i}Z_{j}}$ for each $(i,j)\in E(G_{\theta})$, with time complexity $\mathcal{O}(|E(G)|)=\mathcal{O}(vh^{2}l^{4})$, where $h$ is the frequency of the character appearing the most in $\theta$ except the last string.
  • The mixing Hamiltonian $H_{M}$ requires $\mathcal{O}(kl^{2})$.
Therefore, the total time complexity for quantum operations over $p$ layers is $\mathcal{O}(p(kl^{2}+vh^{2}l^{4})$. We then have the following time complexities for the measurement and classical post-processing:
  • For measuring states we have $\mathcal{O}(kl^{2})$ which we repeat this process $e$ times to get multiple samples. For each sampled bit string, calculating the cost function is $\mathcal{O}\left((kl^{2})^{2}\right)$ which would give us a total time complexity of $\mathcal{O}(e\cdot(kl^{2})^{2})$.
  • For parameter optimization (subgradient method): The classical optimization step updates the parameters $\gamma_{1},\ldots,\gamma_{p}$ and $\beta_{1},\ldots,\beta_{p}$ using a gradient-based method. Each update step takes $\mathcal{O}(p\cdot e)$ for one iteration. Therefore, the total time complexity for parameter optimization is $\mathcal{O}(T\cdot p\cdot e)$.
The last step-by-step analysis gives us the total time complexity of the QAOA algorithm, which is:
$\mathcal{O}(T\cdot p\cdot e\cdot(kl^{2})^{2}+p(kl^{2}+vh^{2}l^{4})).$
This equation shows that Algorithm 2 is a polynomial-time approximate algorithm in all of $T,p,k,l,h$. For higher accuracy, we increase $p$. In practice, $p$ is typically taken to be $\log(\text{no. of vertices})$. This is significantly better than the $2^{kl^{2}/4}$ time complexity of the algorithm if we had used the quantum algorithm proposed by [8]. The implementation of both quantum and classical D0LII solvers developed for this paper is available at: GitHub.
6. Conclusion and Future Directions
In this paper, we introduced the characteristic graph of deterministic L-systems inductive inference. This graph allowed us to map the deterministic L-systems inductive inference problem into maximum independent set problem and the SAT problem. In particular, we proved that the maximum independent set of a certain size in the characteristic graph enables the recovery of a deterministic L-system, which generates the input sequence. Moreover, our development of a quantum algorithm for deterministic L-systems inductive inference is a result that points to another problem in machine learning that benefits from the power of quantum computing. This work not only provided quantum and classical algorithms for deterministic L-systems inductive inference but also opened new avenues for solving other inference problems using graph algorithms. Through our results, one could now translate some of the questions about inference problems of 0L-systems into the language of graph theory. For example, one could easily see that the number of distinct deterministic L-systems systems generating a sequence is precisely equal to the number of distinct maximum independent sets of size $k$ of the characteristic graph of the deterministic L-systems inductive inference, where $k$ is as in Theorem 3.1. Furthermore, it is not hard to see that by removing the first condition of the characteristic graph, one could model an 0L-system inference problem. Another application of this work is that, instead of using QAOA for solving the deterministic L-system inductive inference problem, one may use adiabatic quantum computing, leveraging the fact that we now have a viable candidate for the Hamiltonian for this problem. We leave this for future work.